You are here

Beam delivery update

header

11 December 2014

They cyclotron’s running fairly well, delivering close to 100 µA to BL2C4 for isotope production and protons as needed to BL1B for PIF. BL2A is off at least overnight as we have no RIB users at ISAC – we were able to deliver 211Fr to FrPNC experiment last night, but the experiment suffered a failure shortly thereafter that has forced the group to abandon the remainder of their beam time. The next experiment, collinear laser experiment S1329, isn’t scheduled to begin until Monday. The beamline will be available for tuning tomorrow but the experimenters don’t expect to be ready to begin data-taking until Sunday at the earliest. BL2A will be brought online to allow stable-beam tuning to that experiment sometime before that; a decision as to when will be made tomorrow.

-- ACM